OpenStudy (trwatkins1):

Which list of values is ordered from least to greatest? A. |–5|, 0, 3, –|–4| B. –7, –|–3|, 0, |–4| C. –|6|, |–3|, 1, 4 D. |–12|, –|–5|, 2, |–8|

OpenStudy (trwatkins1):

@SolomonZelman

OpenStudy (trwatkins1):

this is hard

OpenStudy (solomonzelman):

Absolute value of a negative number, will give you the same positive number. For example |-4|=4 (because distance on the number line, from -4 to 0 is same as from 4 to 0, -- 4 units)

OpenStudy (trwatkins1):

yes i know but this is confuseing

OpenStudy (solomonzelman):

Re-write each option and simplify all the terms that have |...|

OpenStudy (rainbow_rocks03):

What do you think the answer is?

OpenStudy (solomonzelman):

like option D for example: \(\large\color{black}{ \displaystyle |–12|,~~~ –|–5|,~~~~ 2, ~~~~|–8| }\) \(\large\color{black}{ \displaystyle 12,~~~ –5,~~~~ 2, ~~~~8 }\) this order is NOT right, because the term |-12| should have been at the very end on the right, not as it is now.

OpenStudy (trwatkins1):

ok

OpenStudy (trwatkins1):

its not a or b

OpenStudy (solomonzelman):

Why not B?

OpenStudy (trwatkins1):

its not in write order

OpenStudy (solomonzelman):

\(|\)-3\(|\)=3 thus, -\(|\)-3\(|\)=-3

OpenStudy (solomonzelman):

and \(|\)-4\(|\)=4

OpenStudy (trwatkins1):